A 10-Newton object moves to the left at 1 m/s. Its kinetic energy is approximately Joules if you double the speed.



a. 0.5 b. 1 c. 10 d. 20



1
Expert's answer
2022-02-28T10:42:33-0500

The kinetic energy


"E_k=\frac{mv^2}{2}=\rm\frac{1\: kg*(1\: m/s)^2}{2}=0.5\:J"

Answer: a